√2(9 √5 - 7 √9)=? I am finding it very hard to solve this type of problem. Can you help me out?
Answer by midwood_trail(310)

√2(9 √5 - 7 √9)=
Use the distributive rule.
Do you remember what that is?
Here is the basic idea:
This means to multiply every number or term that lies inside the parentheses by the number or term that is outside the parentheses.
√2 times (9√5) = 9√10
√2 times (-7√9) = -7√18
We have now removed the parentheses. This is the main reason why we use the distributive rule.
We now have this:
9√10 - 7√18 =
Do you see √18?
We can break this radical down.
√9 times √2 = 3√2
We now have this:
9√10 - 7(3√2) = 9√10 - 21√2
We cannot break this question down anymore.
Final answer: 9√10 - 21√2